OpenStudy (anonymous):

Algebra 1 help please!!! An expression is shown below: f(x) = -16x2 + 60x + 16 Part A: What are the x-intercepts of the graph of the f(x)? Show your work. (2 points) Part B: Is the vertex of the graph of f(x) going to be a maximum or minimum? What are the coordinates of the vertex? Justify your answers and show your work. (3 points) Part C: What are the steps you would use to graph f(x)? Justify that you can use the answers obtained in Part A and Part B to draw the graph. (5 points)

OpenStudy (imstuck):

These are fun! Are you familiar with completing the square to find your vertex and stuff?

OpenStudy (anonymous):

I do remember learning about it but I can't seem to remember at the moment. Can you explain please?

OpenStudy (imstuck):

I will do the problem step by step, explaining along the way what I am doing, ok?

OpenStudy (anonymous):

Alright

OpenStudy (imstuck):

part A: The x intercepts come from factoring the problem and finding where x = 0. Factor that out. BRB with a factored form.

OpenStudy (imstuck):

Ok, the factored form of that gives you the following: -4(x-4)(x+1/4). Setting each of these equal to zero, we have x-4=0 and x+1/4=0. Solving for x we get that x = 4 and x = -1/4. Those are your x intercepts. In other words, those are the points on the x axis where the graph will go through it. Like this, sort of: